# DeCoF: Generated Video Detection via Frame Consistency

### v2

@@ Abstract
The escalating quality of video generated by advanced video generation methods leads to new security challenges in society, which makes generated video detection an urgent research priority. To foster collaborative research in this area, we construct the first open-source dataset explicitly for generated video detection, providing a valuable resource for the community to benchmark and improve detection methodologies. Through a series of carefully designed probe experiments, our study explores the significance of temporal and spatial artifacts in developing general and robust detectors for generated video. Based on the principle of video frame consistency, we introduce a simple yet effective detection model (DeCoF) that eliminates the impact of spatial artifacts during generalizing feature learning. Our extensive experiments demonstrate the efficacy of DeCoF in detecting videos produced by unseen video generation models and confirm its powerful generalization capabilities across several commercial proprietary models.
@@ 1 Introduction
Following the success of diffusion models in generating highly realistic images (Nichol & Dhariwal, 2021; Rombach et al., 2022; Parmar et al., 2023), a growing number of researchers have shifted their focus to the advancement of these models for video generation (Kawar et al., 2023; Blattmann et al., 2023; Zhang et al., 2023). Such developments have turned diffusion models into robust and versatile video generators, enabling the creation of high-quality videos from general conditional inputs, such as text and images. While this powerful and flexible generation capacity has the potential to alleviate the stress of daily life and inspire creativity in work, it concurrently introduces a new crisis of trust (Sharma et al., 2023; Barrett et al., 2023). Therefore, there is an urgent need for detectors capable of identifying generated videos.
In this paper, we propose a challenging task: Generated Video Detection, which aims to create universal detectors to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. To evaluate generated video detectors, we create the first comprehensive generated video detection dataset: the $GeneratedVideoForensics$ (GVF) dataset. The GVF dataset encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments. Moreover, pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos. Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures (Khachatryan et al., 2023; Zhang et al., 2023; Wang et al., 2023a; Sterling., June 2023), such as pixel-space or latent-space models, which aims to measure the generality of the generated video detectors.
Figure 1: Can the detector trained on a specific video generation model cope with the challenges presented by unseen generation models?
Considering that video content contains individual frame contents and smooth consistency across multiple frames, a natural idea is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) models to achieve generated video detection. Our exploratory experiments found that these models have excellent discriminative capabilities for specific T2V models but cannot be applied to detect unseen T2V models.
Refer to caption (a) Geometry
Refer to caption (b) Layout
Refer to caption (c) Frequency
Refer to caption (d) Color
@@
Figure 2:  Spatial and temporal artifacts on generated video. We illustrate the spatial and temporal artifacts present in the generated videos. Spatial artifacts: (a) errors in geometric appearance, (b) errors in image layout, (c) frequency inconsistency, (d) color mismatch; Temporal artifacts: (e) mismatch between frames.
Further probe experiments in Table 1 reveal that the spatiotemporal convolutional models predominantly identify video authenticity by detecting spatial artifacts, resulting in generated video detection that has degenerated into a 2D problem. A critical question arises: Spatial and temporal artifacts, which artifacts are key to constructing universal detectors?
To explore this problem, we first analyze the possibility of building a universal generated video detector through spatial artifacts. We convert the 3D problem of generated video detection into a 2D problem, generated image detection, and implement T2V detection based on existing text-to-image (T2I) detection methods. In section  5.2, we evaluate the performance of these generated image detectors on generated video detection, noting that the reason why these detectors lack generality to unseen generative models is that capturing specific spatial artifacts to identify video authenticity, while spatial artifacts are specific to the T2V model, these detectors are unable to capture more general and high-level spatial artifacts to establish better decision boundaries. So far, we conclude that spatiotemporal convolutional models primarily identify video authenticity by detecting spatial artifacts but are unable to detect more general spatial artifacts. Since spatial artifacts are easily detected, learning of temporal artifacts is difficult to perform.
To rectify this biased learning approach, mitigate the adverse effects induced by spatial artifacts, and construct a more generalizable generated video detector by capturing temporal artifacts, we propose to decouple temporal and spatial artifacts by mapping video frames to feature space, where the distance between features is negatively correlated with image similarity, through a pre-trained CLIP:ViT (Dosovitskiy et al., 2020). More importantly, the features of real video frames are still continuous after being mapped by the mapping function with the above requirements. Meanwhile, the frames of the generated video with inter-frame mismatch will have anomalies in their features. Besides, it greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned. Specifically, we only implement the learning of inter-feature anomalies through a temporal artifact modeling module with two transformer layers (Vaswani et al., 2017). Extensive exploring experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection, especially significant generalization capabilities for unseen generation models.
Our contributions are summarized as follows:
  • We present the first dataset that is comprehensive and scalable as new models emerge for benchmarking generated video detectors.
  • We explore the possibility of detecting generated videos by capturing spatial and temporal artifacts and point out the drawbacks of the detector by capturing spatial artifacts and their causes.
  • We develop a universal generated video detector named DeCof, validating through extensive experiments, including tests on commercial proprietary models.
@@ 2.1 Generation Models
Recently, with the success of diffusion models in generating images (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021; Esser et al., 2021; Rombach et al., 2022; Peebles & Xie, 2023), a number of Research explores text-to-video Diffusion Models (VDMs). Some work (Ramesh et al., 2022; Blattmann et al., 2023; Khachatryan et al., 2023; Wang et al., 2023a)ground their models on latent-based VDMs, for example, Text2Video-zero (Khachatryan et al., 2023) generates videos by using motion dynamics to enhance the latent code of a text-to-image stable diffusion model without requiring training on any video data, and ModelScopeT2V (Wang et al., 2023a) extends it with T2I stable diffusion model of convolutional and attentional blocks and trained using image-text and video-text datasets, ZeroScope (Sterling., June 2023) is a watermark-free T2V generative model trained with the weights of ModelScopeT2V, which is optimized for 16:9 compositions. On the other hand, PYoCo (Ge et al., 2023)and Make-A-Video (Singer et al., 2022) anchor their models on pixel-based VDMs. Show1 (Zhang et al., 2023) integrates pixel-based and latent-based VDMs, using Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, while latent-based VDMs upscale these low-resolution videos from pixel-based VDMs to then create high-resolution videos with low computation cost.
2.2 Generated Content Detections
In recent years, a large amount of work has focused on generated image detection (McCloskey & Albright, 2019; Zhang et al., 2019; Barni et al., 2020; Corvi et al., 2023; Sha et al., 2023), especially for unseen generation models. A simple classifier trained by Wang et al. (Wang et al., 2020) on pagan-generated images using multiple data augmentations can be generalized to 20 unseen other GAN-generated images. Frank et al. (Frank et al., 2020) concluded that the artifacts of the generated images are caused by the upsampling method. Wang et al. (Wang et al., 2023b) identify generated images through the difference of images reconstructed by a pre-trained diffusion model. Tan et al. (Tan et al., 2023) adopt a pre-trained CNN model as transformation models convert images into gradients. Ojha et al. (Ojha et al., 2023) proposed based on pre-trained vision-to-language models to develop a universal generated image detector. In orthogonal, a large amount of work focuses on the detection of Deepfake images (Guarnera et al., 2020; Li et al., 2020; Zhao et al., 2021a; Shiohara & Yamasaki, 2022), such as artificially creating general forged boundaries (Shiohara & Yamasaki, 2022), detect shifts in the frequency domain (Guo et al., 2023), and detect Deepfake from inconsistencies in face image regions (Zhao et al., 2021b).
So far, there are some studies on detecting Deepfake videos(Güera & Delp, 2018; Masi et al., 2020; Agarwal et al., 2020; Haliassos et al., 2021; Wang et al., 2023c), but there is a lack of research specifically focused on detecting generated videos. We hope this paper can bring exploratory and insightful contributions to this domain of study.
@@ 3 Problem Formulation
Generated Video Detection. In addition to the generalization ability of generated video detectors to unseen generation models, we also focus on the generalization to cross-category videos. So, the generated video detection task is defined as follows:
T2V-zero Show1 MScope ZScope Real
Time 0 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: We generated four sub-datasets based on the prompts of real videos and generation models based on four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and SHOW-1. Each pair of positive and negative samples follows the same prompt, and we show the first five frames of four pairs of positive and negative samples corresponding to one of the prompts: A person is cutting a carrot.
Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refers to each real video and its corresponding content prompt. And let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denotes video generation models, where $g_{j}$, $j\in[1,m]$ represents a specific video generation model. Accordingly, based on the prompt $p_{i}$ of each real video $r_{i}$, it generates $m$ corresponding generated videos, resulting in a set of fake video collections $F_{j}\in F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$. the generated video detection dataset $D_{j}$in $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $x\in\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$and $y\in\mathcal{Y}=\{0,1\}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
The goal of generated video detection is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generate model $m_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and achieve generalization on unseen generate models $g{k}$, where $k\in[1,m],k\neq j$. This can be achieved by minimizing classification error of $\mathcal{F}$ on training data $D^{*}_{i}$:
@@
Where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$ We implement the evaluation of detection of cross-category generated videos by ensuring that the training set and the test set do not have videos with the same prompt.
@@ 4 Dataset Construction
To benchmark generated video detectors, we created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ dataset. In this section, we mainly introduce the construction of the data set and conduct a comprehensive evaluation of the dataset.
@@ 4.1 Dataset Composition
The $GeneratedVideoForensics$ (GVF) dataset $\mathcal{D}$ consists of 964 triples and can be represented as $\mathcal{D}=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $F$ represents multiple fake videos generated by different generation models based on the text prompt $P$. In order to examine the generalization of the detectors, we selected four open source generate models with different generation methods and model frameworks: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a), ZeroScope (Sterling., June 2023) and SHOW-1 (Zhang et al., 2023) as generation models for negative samples. Therefore, the GVF dataset could be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show1. Figure  3 shows an example of one of the triples.
Figure 4: Data distribution over categories under the “major content (upper) and “attribute control” (lower) aspects.
Prompt. To approximate the distribution of videos in real task scenarios, we collected 964 prompts and corresponding real videos from existing open-domain text-to-video datasets: MSVD (Chen & Dolan, 2011) and MSR-VTT (Xu et al., 2016) datasets. Following the FTEV (Liu et al., 2023), we consider the collection of prompts from the following four aspects to simulate the sample distribution in real scenarios: the spatial and temporal major content it describes and the spatial and temporal attributes it controls during video generation. Spatial content is classified as people, animals, plants, food, vehicles, buildings, and artifacts, with corresponding spatial attributes of quantity, color, and camera view. Temporal content and attributes encompass actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order. This ensures that our data set is comprehensive and balanced. Figure  4 shows the data distribution on categories under different aspects, which shows that the GVF dataset contains enough data to cover different category ranges.
Figure 5: Manual evaluation results across spatial major contents.
@@ 4.2 Dataset Evaluation
Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation ideas and model architectures, to restore better the challenges posed by invisible generation models. Figure 5 shows the performance of videos generated by the above four generation models on different attributes under the FETV  (Liu et al., 2023) standard. In order to generate different kinds of content, it is obvious to note that each of these models has its advantages and disadvantages. Text2Video-zero, for example, has the best impact on the spatial quality of the video but the worst impact on the temporal continuity. On the contrary, ZeroScope has a better effect on temporal continuity but has the worst effect on spatial quality. The choice of these video generation models, each possessing distinct advantages and disadvantages, ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
5 Methodology
5.1 Probe experiments of STCNN models
Initially, we considered using a spatiotemporal convolutional model (STCNN) to train the generally generated video detector. However, we discovered that these models are only effective on the specific generative models they were trained on and needed more generalizability, failing to discern unseen T2V generation models.
To analyze what information the STCNN network was learning during discrimination, or rather, which features within the video it was utilizing to distinguish between authentic and fabricated content, we designed a sophisticated probe experiment with two novel datasets under the same model framework: The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to see if the model can still differentiate based solely on the temporal discontinuities of the videos. The second dataset was created by randomly selecting a frame from the current test set and replicating it multiple times to see if the model could still differentiate based on the spatial artifacts of the video frame alone.
The experimental results in table 1 led us to the following conclusion: We found that the model’s attention to spatial artifacts and temporal artifacts was distinct, and for developing a generalized discriminator, they did not play the same role. Therefore, we attempt to build detectors based on spatial artifacts and temporal artifacts separately, individually verifying the role each played in the learning process of the universally generated video detector.
Table 1: Spatiotemporal Convolutional Neural Networks (3DCNN) rely on easily detectable spatial artifacts.
@@ 5.1 Probe experiments of STCNN models
5.2 Shortcomings of learning from spatial artifacts
To analyze the role of spatial artifacts, we first work with existing T2I models which have been used as image-level detectors for generated videos, and we test them on the GVF dataset. As shown in Table 2, these detectors demonstrate superior capabilities in capturing the spatial artifacts of specific generation models. However, they lack generalization to unseen generation models.
To find out why these detectors lack generalization to unseen generate models, we visualized the feature space of the detector (Wang et al., 2020) and examined the distribution of four sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, consisting of fake videos from the sub-dataset Text2Video-zero, (iii) Show1, composed of fake videos from the sub-dataset Show-1, and (iv) MP, comprised of fake videos from the sub-dataset ModelScope, (v) ZP, comprised of fake videos from the sub-dataset ZeroScope. For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE (Van der Maaten & Hinton, 2008).
As shown in Figure 6, on the one hand, the detector has ideally classified the real video and the fake video generated by the specific generate model, which means that the detector has fully captured the spatial artifacts of the specific generate model Text2Video-zero  (Khachatryan et al., 2023). On the other hand, the detector cannot distinguish fake videos generated by unseen models from real videos. Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting other video generation models? We illustrate this by visualizing the frequency spectrum of video frames generated by various video generate models. As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent; capturing more general spatial artifacts appears to be extremely difficult.
Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models. The feature space used is one trained to distinguish fake videos (Text2Video-zero) from real videos.
Refer to caption Real
Refer to caption T2V
Refer to caption Show1
Refer to caption MP
Refer to caption ZP

### v3

@@ Abstract
The escalating quality of video generated by advanced video generation methods results in new security challenges, while there have been few relevant research efforts: 1) There is no open-source dataset for generated video detection, 2) No generated video detection method has been proposed so far. To this end, we propose an open-source dataset and a detection method for generated video for the first time. First, we propose a scalable dataset consisting of 964 prompts, covering various forgery targets, scenes, behaviors, and actions, as well as various generation models with different architectures and generation methods, including the most popular commercial models like OpenAI’s Sora and Google’s Veo. Second, we found via probing experiments that spatial artifact-based detectors lack generalizability. Hence, we propose a simple yet effective detection model based on frame consistency (DeCoF), which focuses on temporal artifacts by eliminating the impact of spatial artifacts during feature learning. Extensive experiments demonstrate the efficacy of DeCoF in detecting videos generated by unseen video generation models and confirm its powerful generalizability across several commercially proprietary models. Our code and dataset will be released at https://anonymous.4open.science/r/DeCoF-8394.
@@ 1 Introduction
Following the success of diffusion models in generating highly realistic images [28, 34, 30], increasing attention has been paid to video generation [21, 4, 51], enabling the creation of high-quality videos from general conditional inputs such as text and images. However, this powerful generation capacity also introduces a new crisis of trust [36, 3]. Therefore, there is an urgent need for detectors for generated videos.
Figure 1: Can the detector trained on a specific video generation model detect videos generated by unseen generation models?
In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. Moreover, we propose the first comprehensive generated video detection dataset, the $GeneratedVideoForensics$ (GVF) dataset, for evaluating the performances of generated video detectors. GVF encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments. Pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos. Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures, such as pixel-space or latent-space models [22, 51, 46, 42], to measure the generalizability of the generated video detectors.
(a) Geometry
(b) Layout
(c) Frequency
(d) Color
@@
Figure 2: Illustration of spatial and temporal artifacts on generated video. Spatial artifacts: (a) errors in geometric appearance, (b) errors in image layout, (c) frequency inconsistency, (d) color mismatch; Temporal artifacts: (e) mismatch between frames.
Considering that authentic videos should present smooth inter-frame consistency, a naive approach is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) for generated video detection. However, we found via exploratory experiments that these models lack generalizability to detect videos generated by unseen T2V models.
Further probe experiments revealed that spatiotemporal convolutional models primarily recognize video authenticity through spatial artifacts (Figure 2 illustrates two types of artifacts in generated videos), degrading video detection into a 2D problem. By analyzing the performance of existing text-to-image (T2I) detection works [48, 43, 47], we found it difficult to capture general spatial artifacts, as the contrast and style of video frames generated by different generation models vary.
Therefore, to develop a universal generated video detector, we propose the first temporal-artifact-based detection method. More specifically, we decouple temporal and spatial artifacts by mapping video frames to feature space where the inter-feature distance is negatively correlated with image similarity, to detect the anomalies caused by inter-frame inconsistency. Our method greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned. Extensive experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection and excellent generalizability for unseen generation models, even on the latest generation models Sora [41] and Veo [45].
Our contributions can be summarized as follows:
  • We present the first comprehensive and scalable dataset for benchmarking generated video detectors.
  • Through a series of probing experiments, we have demonstrated the flaws in detecting generated videos from spatial artifacts.
  • We develop the first temporal-artifact-based universal generated video detector named DeCof and validate its performance through extensive experiments, including tests on the latest commercial proprietary models.
@@ 2.1 Generation Models
Recently, with the success of diffusion models in generating images [20, 40, 28, 9, 34, 31], a number of research explores text-to-video Diffusion Models (VDMs). Some works [33, 4, 22, 46] ground their models on latent-based VDMs, e.g., Text2Video-zero [22] generates videos by using motion dynamics to enhance the latent code of text-to-image(T2I) model [34] without the need for training on any video data, and ModelScopeT2V [46] extends T2I model with temporal convolution and attention blocks trained with image-text and video-text datasets. ZeroScope [42] is a watermark free version of ModelScopeT2V, and has been optimized for 16:9 composition. On the other hand, PYoCo [14] and Make-A-Video [38] anchor their models on pixel-based VDMs. Show1 [51] integrates Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, then upscale them with latent-based VDMs to create high-resolution videos with low computation cost.
T2V-zero
Show1
MScope
ZScope
Real
Time 0
Time 1
Time 2
Time 3
Time 4
Figure 3: We generated four sub-datasets based on the real videos using the prompts and selected generation models employing four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1. Each pair of positive and negative samples use the same prompt. Here shows the first five frames of four sample pairs using the prompt: A person is cutting a carrot.
2.2 Generated Content Detection
Many studies have focused on generated image detection [27, 52, 2, 7, 35], especially for unseen generation models. A simple classifier trained by Wang et al. [47] on ProGAN-generated images can be generalized to 20 unseen other GAN-generated images. Frank et al. [13] conclude that the artifacts of the generated images are caused by the upsampling method. Wang et al. [48] identify generated images through the difference of images reconstructed by a pre-trained diffusion model. Ojha et al. [29] propose to develop a universal generated image detector based on pre-trained vision-to-language models. Meanwhile, a large amount of work focuses on Deepfake image detection [16, 23, 53, 37], using methods such as artificially creating general forged boundaries [37], detecting shifts in the frequency domain [18], and detecting inconsistencies in face image regions [54]. So far, there are some studies on detecting Deepfake videos [17, 26, 1, 19, 49], but research on detecting generated videos remain unexplored. This paper targets at bringing exploratory and leverageable contributions to this new direction.
@@ 3 Problem Formulation
Generated Video Detection. In addition to the generalizability to unseen generation models, we also take the generalizability to cross-category videos into account. The generated video detection task can be defined as follows:
Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refer to each real video and its corresponding content prompt. Let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denote a set of video generation models. Using the prompt $p_{i}$ of each real video $r_{i}$, $G$ generates $m$ corresponding videos, resulting in a set of fake video collections $F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$, $j\in[1,m]$. As such, we form the generated video detection dataset $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$, $\mathcal{Y}=\{0,1\}$, $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
The goal is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generation model $g_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and generalizes to unseen generation models $G_{k}$, where $k\in[1,m],k\neq j$. This can be achieved by minimizing classification error of $\mathcal{F}$ on training data $D^{*}_{i}$:
@@
where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$. We evaluate the detection of cross-category generated videos by ensuring that the training set and the test set do not contain videos using the same prompt.
@@ 4 Dataset Construction
To benchmark generated video detectors, we have created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ (GVF) dataset.
@@ 4.1 Dataset Composition
The GVF dataset consists of 964 triples and can be represented as $\mathcal{D}=\{(r,p,F)^{(i)}\}_{i=1}^{964}$, where $r$ and $p$ represent the real video and the corresponding text prompt, and $F$ represents the videos generated by the generation models using the text prompt $p$. In order to examine the generalizability of the detectors, we have selected four open source generation models employing different generation methods and model frameworks, namely Text2Video-zero [22], ModelScopeT2V [46], ZeroScope [42] and Show-1 [51] as generation models for negative samples. Accordingly, the GVF dataset could be divided into four sub-datasets. Figure 3 shows an example triple.
Figure 4: Data distribution across categories under the “major content (upper) and “attribute control" (lower) aspects.
Figure 5: Performance comparison of the 4 selected models across spatial major contents.
Prompt. To approximate the distribution of videos in reality,we have collected 964 prompts and corresponding real videos from open-domain text-to-video datasets: MSVD [6] and MSR-VTT [50]. Following FTEV [25], we consider the collection of prompts from the following four aspects to simulate the sample distribution in reality: the spatial and temporal content it mainly describes and the spatial and temporal attributes it controls during video generation. Spatial content is classified as people, animals, plants, food, vehicles, buildings, and artifacts, with corresponding spatial attributes of quantity, color, and camera view. Temporal content encompasses actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order. This ensures that our GVF dataset is comprehensive and balanced. Figure 5 illustrates the data distribution across categories under different aspects, which indicates that the GVF dataset contains enough data to cover different category ranges.
@@ 4.2 Dataset Evaluation
Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation methods and model architectures, to simulate the challenges posed by unseen generation models. Figure 5 shows the generation ability of the selected generation models on different attributes under the FETV [25] standard. Clearly, each model has its advantages and disadvantages. Text2Video-zero, for example, has the best performance on the spatial quality of the video, but the worst performance on the temporal continuity. On the contrary, ZeroScope has a better performance on temporal continuity but the worst performance on spatial quality. The choice of these video generation models ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
Table 1: Spatiotemporal Convolutional Neural Networks (STCNN) rely on spatial artifacts.
@@
5 Methodology
5.1 Probe experiments of STCNN models
Initially, we considered using spatiotemporal convolutional model(STCNN) [5, 12, 11, 10], which are widely used in video classification tasks, to train the generalizable generated video detector. However, we discovered that these models are only effective on the specific generation models they were trained on, failing to discern unseen T2V generation models.
To analyze what information the STCNN network was learning during the training process, or rather, which features within the video it utilized to distinguish between authentic and fabricated content, we designed a probing experiment with two novel datasets under the same model framework. The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to observe if the model can differentiate based solely on the temporal discontinuities of the videos. The second dataset is created by randomly selecting a frame of fake videos from the current test set and replicating it multiple times to observe if the model could differentiate based on the spatial artifacts of the video frame alone.
The experiment results in Table 1 show that STCNN overly relies on spatial artifacts, degrading the generated video detection into a 2D problem. However, the fact STCNN relies on spatial artifacts still does not explain its lack of generalizability to unseen generation models.
To further explore the possibility of detecting generated video based on spatial artifacts, we turned to the SoTA generated image detection works.
5.2 Capturing Universal Spatial Artifacts is Difficult
Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models.
We transform the current detection of generated images into image level detectors for detecting generated videos and test them on the GVF dataset. As shown in Table 2, these detectors demonstrate superior capabilities in capturing the spatial artifacts of specific generation models. However, they also lack generalizability to unseen generation models.
To find out why these detectors lack generalizability to unseen generate models, we visualized the feature space of the detector [47] and examined the distribution of five sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, comprising fake videos from the sub-dataset Text2Video-zero, (iii) Show1, comprising fake videos from the sub-dataset Show-1, and (iv) MP, comprising fake videos from the sub-dataset ModelScope, (v) ZP, comprising fake videos from the sub-dataset ZeroScope.
For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE [44].
Real
T2V
Show1
MP
ZP